# TAMS: Translation-Assisted Morphological Segmentation

### v1

@@ Abstract
Canonical morphological segmentation is the process of analyzing words into the standard (aka underlying) forms of their constituent morphemes. This is a core task in language documentation, and NLP systems have the potential to dramatically speed up this process. But in typical language documentation settings, training data for canonical morpheme segmentation is scarce, making it difficult to train high quality models. However, translation data is often much more abundant, and, in this work, we present a method that attempts to leverage this data in the canonical segmentation task.
We propose a character-level sequence-to-sequence model that incorporates representations of translations obtained from pretrained high-resource monolingual language models as an additional signal. Our model outperforms the baseline in a super-low resource setting but yields mixed results on training splits with more data. While further work is needed to make translations useful in higher-resource settings, our model shows promise in severely resource-constrained settings.
@@ 1 Introduction
Morphological segmentation is the task of breaking words into morphemes, the smallest semantic units of a language. Morphemes can merge and change during word formation, and the precise morphological composition of a word is often obfuscated in its surface form. Segmentation can thus take two forms: surface/linear segmentation and canonical segmentation, which divides a word into the “canonical” forms of its morphemes (cf. Figure 1).
One important motivation for automated canonical segmentation is to expedite the process of linguistic analysis, including the creation of Interlinear Glossed Text (IGT). IGT is a form of morphological annotation that typically adheres to the Leipzig glossing format Lehmann (1982), a documentation style wherein each line of the target text is broken up into a transcription line, a gloss line (morphological annotation), and a translation line. IGT is a crucial resource in language documentation work, but it is costly and time-consuming to generate. The task of morphological segmentation is a key component in glossing, and automated canonical segmentation could aid in this process. Prior work has shown that automated methods have the potential to assist language documentation and revitalization Palmer et al. (2009); Moeller et al. (2020); Moeller and Hulden (2021); Chaudhary et al. (2022); Ahumada et al. (2022).
@@
Neural models have been shown to perform well on the task of canonical segmentation (Kann et al., 2016; Ruzsics and Samardžić, 2017), but the success of these models has been restricted by the availability of annotated segmentation data. IGT, though a limited resource, is one important source of training data for canonical segmentation. However, until now, primarily the transcription and gloss lines of IGT have been used as input to segmentation models, while the translations have been overlooked. Moreover, in real-world documentation settings, translated data is often much more available than morphologically analyzed data, making it practically attractive as an additional input. Here we consider how supervised canonical segmentation methods can be improved by leveraging this underutilized data source. Translations also provide the opportunity to make use of pretrained models that likely have much higher-quality representations than any model for the target language.
@@
Our work is inspired in part by Zhao et al. (2020) who experiment with leveraging translations for the task of automatic interlinear glossing. They use a multi-source word-level transformer to jointly model the transcription and translation sequences, and outperform previous baselines. Their work shows promise for the utility of translation data in morphological analysis tasks. However, they assume the presence of well-segmented data and state that "proper segmentation remains a challenge and that the creation of segmentation tools is a valuable endeavor." Our work endeavors to address the segmentation issue.
@@
We treat canonical segmentation as a sequence-to-sequence problem and use a character-level pointer-generator LSTM See et al. (2017) to map each surface form word to its segmented, canonicalized form. We experiment with Tsez and Lezgi, two Northeast Caucasian languages, and Arapaho, a Plains Algonquian language. We leverage existing sentence-level English translations present in the IGT data from the SIGMORPHON 2023 Shared Task on Interlinear Glossing Ginn et al. (2023) and create two datasets of word-level transcription–translation alignments: automatically with awesome-align [^1] Licensed under the BSD 3-Clause License.Dou and Neubig (2021), and manually according to conventions described in §4.3.2. We then embed these translations with BERT Devlin et al. (2019) and experiment with incorporating them into our baseline model’s encoder and decoder. We also analyze the impact of training set size on the efficacy of our approach by limiting our training split to simulate varied levels of resourcedness. Finally, we analyze the effect of automatic vs. manual word-alignments on the subset of the Tsez data that we manually aligned. [^2] We will make our manual alignments publicly available upon acceptance.
We find that our approach is most impactful in the extremely low data setting (n=100) and outperforms the baseline by an average of 2.33% and as high as 3.88% accuracy. However, in the higher data settings, incorporating translations may or may not be beneficial. Our results also suggest that gold aligned data may not be necessary to see improvements over the baseline.
@@ 2 Related Work
Previous work has suggested that distributional similarity is an informative cue for morphology (Yarowsky and Wicentowski, 2000; Schone and Jurafsky, 2001) and static word embeddings encode some morphological information (Musil, 2021; Soricut and Och, 2015). Other work has suggested that BERT embeddings could encode grammatical and morphological information Nastase and Merlo (2023); Jawahar et al. (2019). BERT embeddings have also been used for part-of-speech tagging Tsai et al. (2019); Singh et al. (2021); Mohseni and Tebbifakhr (2019). We aim to leverage the morphological cues intrinsic to pretrained embeddings of English translations to improve our segmentation models.
@@ 3 Incorporating Translations into Morphological Segmentation Models
Here, we present our method for using translations as a source of additional signal for morphological analysis. Our method relies on alignments to identify word forms in the translation that are relevant for the target word, and embeddings of words in the translation obtained from a high-resource language model are used to assist in the segmentation of the word. We describe several approaches for turning these embeddings into a fixed-length representation and for incorporating them as input to the segmentation model.
@@ 3.2 Translation Assistance
The translations of textual data from low-resource languages are usually written in a high-resource language such as English or Spanish. High-resource languages have very high-quality pretrained language models (PLMs) and therefore rich word representations available to them, and we hypothesize that incorporating this signal into the process of morphological segmentation may be helpful. For example, information from the high-resource language might help the segmenter resolve lexical ambiguities. Here, we propose several related methods for incorporating information from a high-resource translation into an RNN-based encoder–decoder morphological segmenter. For clarity, we will concretely consider a unidirectional LSTM-based encoder, though our approach is trivially applicable to other RNN-based encoder–decoder architectures. We refer to the network’s embedding and hidden dimension sizes as $\mathrm{emb}$ and $\mathrm{hid}$, respectively.
@@
A preliminary step is to produce an word-level alignment between source words and translation words like so, where $\mathrm{align}$ represents an aligner’s decision on whether two words are aligned:
@@
For each word $\mathbf{x}$, we now have some aligned translation word representations $\mathbf{d}_{\mathrm{align}}=d_{a},\ldots,d_{b}$. We next produce $v$, a fixed-length representation of $\mathbf{d}_{\mathrm{align}}$ which will be of length $\mathrm{emb}$. We investigate three different strategies for producing this representation which differ in how they treat the sentence-wide representation $d_{0}$.
For the CLS-None strategy, we discard $d_{0}$ and average pool $\mathbf{d}_{\mathrm{align}}$ before using a model parameter $W_{\mathrm{trans}}\in\mathbb{R}^{\mathrm{h}_{\mathrm{PLM}}\times\mathrm{emb}}$ to project the vector from the hidden size of the PLM to the embedding size of the model:
@@
After we have computed $v$, we must somehow incorporate it into the encoder and/or the decoder’s process.
For Concat, we double the model’s input size to $2\times\mathrm{emb}$ and concatenate $v$ to the input embedding at each time step in the LSTM. Concat-Half is identical to Concat, except the model’s input size is held constant, with character embeddings and $v$ sharing the dimensions equally. The model’s character embedding module and $W_{\mathrm{trans}}$ above have their output dimensions halved accordingly. For Init-State, assuming that there is some integer $z$ such that $z\times\mathrm{emb}=\mathrm{hid}$, we initialize the LSTM’s hidden state as $z$ concatenations of $v$ to itself, $\bigoplus_{1}^{z}v$. For Init-Char, we modify the LSTM’s input sequence so that $v$ appears first, as if it were the embedding of a character. All strategies are applicable to either the encoder or the decoder; we experiment with most combinations, except for Init-Char in the decoder.
@@ 4 Data
To perform our experiments, we use IGT data from the SIGMORPHON 2023 Shared Task on Interlinear Glossing Ginn et al. (2023) in three languages: Lezgi, Tsez, and Arapaho. All of the data is licensed under CC BY-NC 4.0. Each word in this IGT format has a surface form, a canonical form, morpheme-level glosses, and an English translation for the sentence the word appears in. An example is shown in Figure 2, which also provides a sample of our manual alignment process.
@@ 4.1.3 Arapaho
Arapaho [arp] is an Algonquian language that is highly agglutinating and polysynthetic. Noun stems can be inflected for plurality, obviation, vocative, and locative cases through suffixing. Nouns also necessarily belong to either animate or inanimate gender, and gender impacts the surface representation of many inflectional markers. Arapaho nouns also participate in derivational morphology, and modified nouns can be derived from independent nouns or verbs. Arapaho verbal morphology is even more complex. In terms of inflectional morphology, verb stems can be divided into four different classes that each take different markers for person, number, and obviation. Verbs can also be broken up into four different orders– affirmative, non-affirmative, conjunct, and imperative– which also impact the inflectional morphemes. Arapaho derivational verbal morphology is extensive, and unique verb forms can be derived through processes of prefixation, suffixation, denominalization, reduplication, and noun incorporation. We consider Arapaho to be an out-of-family test language.
@@ 4.2 Preprocessing
The transcription and translation lines are not always pretokenized in these datasets (e.g., punctuation sometimes appears next to words) and it is necessary to tokenize the data for this reason. We use HuggingFace transformers’ Wolf_Debut_Sanh_Chaumond_Delangue_Moi_Cistac_Rault_Louf_Funtowicz_et BertPreTokenize pretokenizer for this purpose, with some additional processing to make language-specific corrections. (In Arapaho, for example, the apostrophe character ' represents a consonant, and it should not be separated from words it appears in.) After this processing, we verify that there are equal numbers of surface and canonical forms, and we discard any sentences for which this is not true. Finally, we initialize our training instances by finding all unique pairs of surface and canonical forms at the word level and choose one randomly if there is more than one occurrence of it in the corpus. Both surface and canonical forms are NFD normalized. Our full datasets consist of 53800 words in Arapaho, 10952 words in Tsez, and 2060 words in Lezgi. In our experiments, the Arapaho dataset is downsampled to 16666 words in order to bring its size closer to the other two datasets.
@@ 5 Experiment
We experiment with the strategies described in §3.2 for incorporating information from translations into our morphological segmentation model. We proceed by first tuning our exact approach on the development split of a single language, Tsez, by exhaustively considering every combination of encoder, decoder, and CLS-token translation incorporation strategies. We then apply our top-performing model to the test splits of all three languages. All experiments are performed on NVIDIA A100 GPUs, and all model implementations were based on those provided by Yoyodyne. [^3] https://github.com/CUNY-CL/yoyodyne
@@ 5.3 Model and Hyperparameters
The pointer-generator LSTM differs from a regular LSTM encoder-decoder in that it has a pointer network Vinyals et al. (2015), which allows the model to copy over specific characters in the input sequence to the output sequence.
The decoder assesses the probability of copying an element from the input to the output rather than generating it, then computes the probability distribution of the output at each time step by combining the probability distribution across the output vocabulary with the attention distribution over the input characters. The weights, indicating the probability of generation or copying, are determined by a feedforward network.
@@
The results of hyperparameter tuning are very similar for the 500-sentence and full data settings, but vary notably for the 100-sample setting.
In all subsequent experiments, we train our 100 training sample models with one set of hyperparameters and all other models with a separate set. In the 100-sample setting, we use a batch size of 16, two encoder and two decoder layers, an embedding size of 512 and a hidden size of 1024. We set dropout to $3.662\times 10^{-1}$, learning rate to $2.411\times 10^{-4}$, and train for up to 607 epochs.
@@
Both settings use the Adam optimizer and the ReduceLROnPlateau scheduler.
Incorporating Translation Information

### v2

@@ Abstract
Canonical morphological segmentation is the process of analyzing words into the standard (aka underlying) forms of their constituent morphemes. This is a core task in endangered language documentation, and NLP systems have the potential to dramatically speed up this process. In typical language documentation settings, training data for canonical morpheme segmentation is scarce, making it difficult to train high quality models. However, translation data is often much more abundant, and, in this work, we present a method that attempts to leverage translation data in the canonical segmentation task. We propose a character-level sequence-to-sequence model that incorporates representations of translations obtained from pretrained high-resource monolingual language models as an additional signal. Our model outperforms the baseline in a super-low resource setting but yields mixed results on training splits with more data. Additionally, we find that we can achieve strong performance even without needing difficult-to-obtain word level alignments. While further work is needed to make translations useful in higher-resource settings, our model shows promise in severely resource-constrained settings.
@@ 1 Introduction
Morphological segmentation is the task of breaking words into morphemes, the smallest semantic units of a language. Morphemes can merge and change during word formation, and the precise morphological composition of a word is often obfuscated in its surface form. Segmentation can thus take two forms: surface/linear segmentation and canonical segmentation, which divides a word into the “canonical” forms of its morphemes (cf. Figure 1). One important motivation for automated canonical segmentation is to expedite the process of linguistic analysis, including the creation of Interlinear Glossed Text (IGT). IGT is a form of morphological annotation that typically adheres to the Leipzig glossing format Lehmann (1982), a linguistic representation wherein each line of the target text is broken up into a transcription line, a gloss line (morphological annotation), and a translation line. IGT is a crucial resource in endangered language documentation work, but it is costly and time-consuming to generate. The task of morphological segmentation is a key component in glossing, and automated canonical segmentation could aid in this process. Prior work has shown that automated methods have the potential to assist language documentation and revitalization Palmer et al. (2009); Moeller et al. (2020); Moeller and Hulden (2021); Chaudhary et al. (2022); Ahumada et al. (2022).
@@
Neural models have been shown to perform well on the task of canonical segmentation (Kann et al., 2016; Ruzsics and Samardžić, 2017), but the success of these models has been restricted by the availability of annotated segmentation data. IGT, though a limited resource, is one important source of training data for canonical segmentation. Until now, primarily the transcription and gloss lines of IGT have been used as input to segmentation models, while the translations have been overlooked. Moreover, in real-world language documentation settings, translated data is often much more available than morphologically analyzed data, making it practically attractive as an additional input. A typical language documentation pipeline begins with transcription, followed by translation, followed only then by morphological analysis. This commonly results in a setting where a linguist will have a wealth of translation data but comparatively little morphological segmentation data. Here we consider how supervised canonical segmentation methods can be improved by leveraging this underutilized supplemental data source. Also, because they are typically into more widely-spoken languages, translations provide the opportunity to make use of pretrained models that likely have much higher-quality representations than any model available for the low-resource target language.
@@
Our work is inspired in part by Zhao et al. (2020) who experiment with leveraging translations for the task of automatic interlinear glossing (i.e., generating the gloss line of IGT). They use a multi-source word-level transformer to jointly model the transcription and translation sequences, and outperform previous baselines. Their work shows promise for the utility of translation data in morphological analysis tasks. However, they assume the presence of well-segmented data and state that "proper segmentation remains a challenge and that the creation of segmentation tools is a valuable endeavor." Our work endeavors to address the segmentation issue.
@@
We treat canonical segmentation as a sequence-to-sequence problem and use a character-level pointer-generator LSTM See et al. (2017) to map each surface form word to its segmented, canonicalized form. We experiment with Tsez and Lezgi, two Northeast Caucasian languages, and Arapaho, a Plains Algonquian language. We leverage existing sentence-level English translations present in the IGT data from the SIGMORPHON 2023 Shared Task on Interlinear Glossing Ginn et al. (2023) and create two datasets of word-level transcription–translation alignments: automatically with awesome-align [^1] Licensed under the BSD 3-Clause License.Dou and Neubig (2021), and manually according to conventions described in §4.3.2. We then embed these translations with BERT Devlin et al. (2019) and experiment with incorporating them into our baseline model’s encoder and decoder. We also analyze the impact of training set size on the efficacy of our approach by limiting our training split to simulate varied levels of resourcedness. Finally, we analyze the effect of automatic vs. manual word-alignments on a subset of the Tsez data.  [^2] https://github.com/lgessler/tama/blob/master/data/tsez/wat/Tsez%20Train%20101-310.json Based on poor automatic alignment performance, we introduce an additional model variant (TAMS-CLS) which incorporates the translation only at the sentence level.
We find that although gold alignment does lead to better performance with our approach, we still see improvement over the baseline with automatic alignments in some cases. In the extremely low data setting (n=100), our approach outperforms the baseline by an average of 1.99 percentage points and as high as 2.87 points, even with poor quality automatic alignments. In many cases, across training set sizes, TAMS-CLS is the highest performing model configuration, which suggests that we may not even need word-level alignments to achieve performance improvements with TAMS. These results are promising in the context of the extremely resource constrained language documentation setting. However, in higher data settings, incorporating translations may or may not be beneficial. Our results also suggest that aligning translations may not be necessary to see improvements over the baseline.
@@ 2 Related Work
Previous work has suggested that distributional similarity is an informative cue for morphology (Yarowsky and Wicentowski, 2000; Schone and Jurafsky, 2001), and that static word embeddings encode some morphological information (Musil, 2021; Soricut and Och, 2015). Other work has suggested that BERT embeddings could encode grammatical and morphological information Nastase and Merlo (2023); Jawahar et al. (2019). BERT embeddings have also been used for part-of-speech tagging Tsai et al. (2019); Singh et al. (2021); Mohseni and Tebbifakhr (2019). We aim to leverage the morphological cues intrinsic to pretrained embeddings of English translations to improve our segmentation models.
@@ 3 Incorporating Translations into Morphological Segmentation Models
Here, we present our method for using translations as a source of additional signal for morphological analysis. We first align the word forms in our target language to relevant word forms in the translation. We then obtain embeddings of these word-forms from a high-resource language model. We describe several approaches for turning these embeddings into a fixed-length representation and for incorporating them as input to the segmentation model.
@@ 3.2 Translation Assistance
The translations of textual data from low-resource languages are usually written in a high-resource language such as English or Spanish. High-resource languages have very high-quality pretrained language models (PLMs) and therefore rich word representations available to them, and we hypothesize that incorporating this signal into the process of morphological segmentation may be helpful. For example, information from the high-resource language might help the segmenter resolve lexical ambiguities.
Here, we propose several related methods for incorporating information from a high-resource translation into an RNN-based encoder–decoder morphological segmenter. For clarity, we will concretely consider a unidirectional LSTM-based encoder, though our approach is trivially applicable to other RNN-based encoder–decoder architectures. We refer to the network’s embedding and hidden dimension sizes as $\mathrm{emb}$ and $\mathrm{hid}$, respectively.
@@
A preliminary step is to produce alignments between source and translation words like so, where $\mathrm{align}$ represents an aligner’s decision on whether two words are aligned:
@@
For each word $\mathbf{x}$, we now have some aligned translation word representations $\mathbf{d}_{\mathrm{align}}=d_{a},\ldots,d_{b}$. We next produce $v$, a fixed-length representation of $\mathbf{d}_{\mathrm{align}}$ which will be of length $\mathrm{emb}$. We investigate three different strategies for producing this representation which differ in how they treat the sentence-wide representation $d_{0}$. The intuition behind including the CLS token in our representation is that it may allow us to capture sentence-level dynamics better than word-level alignments alone. For the CLS-None strategy, we discard $d_{0}$ and average pool $\mathbf{d}_{\mathrm{align}}$ before using a model parameter $W_{\mathrm{trans}}\in\mathbb{R}^{\mathrm{h}_{\mathrm{PLM}}\times\mathrm{emb}}$ to project the vector from the hidden size of the PLM to the embedding size of the model:
@@
After we have computed $v$, we need to incorporate it into the encoder and/or the decoder’s process. We consider four different strategies for incorporation, described next.
For Concat, we double the model’s input size to $2\times\mathrm{emb}$ and concatenate $v$ to the input embedding at each time step in the LSTM. Concat-Half is identical to Concat, except the model’s input size is held constant, with character embeddings and $v$ sharing the dimensions equally. The model’s character embedding module and $W_{\mathrm{trans}}$ above have their output dimensions halved accordingly. For Init-State, assuming that there is some integer $z$ such that $z\times\mathrm{emb}=\mathrm{hid}$, we initialize the LSTM’s hidden state as $z$ concatenations of $v$ to itself, $\bigoplus_{1}^{z}v$. For Init-Char, we modify the LSTM’s input sequence so that $v$ appears first, as if it were the embedding of a character.
All strategies are applicable to either the encoder or the decoder; we experiment with most combinations, except for Init-Char in the decoder.
@@ 4 Data
To perform our experiments, we use IGT data from the SIGMORPHON 2023 Shared Task on Interlinear Glossing Ginn et al. (2023) in three languages: Lezgi, Tsez, and Arapaho. All of the data is licensed under CC BY-NC 4.0. Each word in this IGT format has a surface form, a canonical segmentation, morpheme-level glosses, and an English translation for the sentence the word appears in. An example is shown in Figure 2, which also provides a sample of our manual alignment process.
@@ 4.1.3 Arapaho
Arapaho [arp] is an Algonquian language that is highly agglutinating and polysynthetic. Noun stems can be inflected for plurality, obviation, vocative, and locative cases through suffixing. Nouns also necessarily belong to either animate or inanimate gender, and gender impacts the surface representation of many inflectional markers. Arapaho nouns also participate in derivational morphology, and modified nouns can be derived from independent nouns or verbs.
Arapaho verbal morphology is even more complex. In terms of inflectional morphology, verb stems can be divided into four different classes that each take different markers for person, number, and obviation. Verbs can also be broken up into four different orders– affirmative, non-affirmative, conjunct, and imperative– which also impact the inflectional morphemes. Arapaho derivational verbal morphology is extensive, and unique verb forms can be derived through processes of prefixation, suffixation, denominalization, reduplication, and noun incorporation. We consider Arapaho to be particularly interesting as an out-of-family test language because of its rich morphology that is notably distinct from that of Tsez.
@@ 4.2 Preprocessing
The transcription and translation lines are not always pretokenized in these datasets (e.g., punctuation sometimes appears next to words), and it is necessary to tokenize the data for this reason. We use HuggingFace transformers’ Wolf et al. (2020) BertPreTokenize pretokenizer for this purpose, with some additional processing to make language-specific corrections. (In Arapaho, for example, the apostrophe character ' represents a consonant, and it should not be separated from words it appears in.) After processing, we verify that there are equal numbers of surface and canonical forms, and we discard any sentences for which this is not true. Finally, we initialize our training instances by finding all unique pairs of surface and canonical forms at the word level and choose one randomly if there is more than one occurrence of it in the corpus. Both surface and canonical forms are NFD normalized. Our full datasets consist of 53800 words in Arapaho, 10952 words in Tsez, and 2060 words in Lezgi. In our experiments, the Arapaho dataset is downsampled to 16666 words, to bring its size closer to the other two datasets.
@@ 5 Experiment
We experiment with the strategies described in §3.2 for incorporating information from translations into our morphological segmentation model. We proceed by first tuning our exact approach on the development split of a single language, Tsez, by exhaustively considering every combination of encoder, decoder, and CLS token translation incorporation strategies. We then apply our top-performing model to the test splits of all three languages. All experiments are performed on NVIDIA A100 GPUs, and all model implementations were based on those provided by Yoyodyne. [^3] https://github.com/CUNY-CL/yoyodyne
@@ 5.2 Evaluation Metrics
Tsez Lezgi Arapaho Average
Train Metrics TAMS TAMS- Base TAMS TAMS- Base TAMS TAMS- Base TAMS TAMS- Base
Limit CLS CLS CLS CLS
Acc. 24.78 23.83 22.19 29.71 29.71 26.84 15.05 15.18 14.00 23.18 22.91 21.01
n = F1 48.37 48.07 47.52 40.62 41.23 38.33 37.48 37.00 36.25 42.16 42.10 40.70
100 ED 3642 3701 3805 793 773 861 30759 30906 32742 11731 11793 12469
Acc. 34.52 34.01 33.36 34.80 35.63 35.19 20.66 21.58 21.50 29.99 30.41 30.02
n = F1 58.13 58.24 57.83 50.24 51.54 51.36 44.69 45.35 45.85 51.02 51.04 51.68
250 ED 3061 3055 3286 772 706 716 28561 27550 29048 10798 10437 11017
Acc. 47.67 47.69 47.91 41.41 41.26 41.60 33.84 34.32 33.70 40.97 41.09 41.07
n = F1 68.77 68.65 68.80 57.64 56.64 57.28 56.48 56.97 56.54 60.96 60.75 60.87
500 ED 2213 2195 2217 617 633 647 19743 19790 20211 7524 7539 7692
Acc. 80.78 81.96 82.60 46.84 47.09 44.66 67.72 67.40 67.08 65.11 65.48 64.78
n = F1 89.52 90.08 90.44 62.48 62.48 60.75 81.62 81.45 81.11 77.87 78.00 77.43
all ED 701 643 652 532 537 568 9899 9970 10495 3711 3717 3905
Table 1: Final results per language: Performance on all languages’ test sets (averaged over 5 randomized test sets) on silver-aligned data. Metrics: Accuracy (Acc.), F1 score (F1), and Edit Distance (ED).
@@ 5.3 Model and Hyperparameters
The pointer-generator LSTM differs from a regular LSTM encoder-decoder in that it has a pointer network Vinyals et al. (2015), which allows the model to copy over specific characters in the input sequence to the output sequence. The decoder assesses the probability of copying an element from the input to the output rather than generating it, then computes the probability distribution of the output at each time step by combining the probability distribution across the output vocabulary with the attention distribution over the input characters. The weights, indicating the probability of generation or copying, are determined by a feedforward network.
@@
The results of hyperparameter tuning are very similar for the 500-word and full data settings, but vary notably for the 100-sample setting. In all subsequent experiments, we train our 100 training sample models with one set of hyperparameters and all other models with a separate set.
In the 100-sample setting, we use a batch size of 16, two encoder and two decoder layers, an embedding size of 512 and a hidden size of 1024. We set dropout to $3.662\times 10^{-1}$, learning rate to $2.411\times 10^{-4}$, and train for up to 607 epochs.
@@
Both settings use the Adam optimizer and the ReduceLROnPlateau scheduler. [^4] https://pytorch.org/docs/stable/generated/torch.optim.lr_scheduler.ReduceLROnPlateau.html.
5.3.1 Alignment Considerations
Precision Recall F1
0.1637 0.1846 0.1735
Table 2: Performance of awesome-align judged against 500 manually aligned Tsez words
Unfortunately, awesome-align does not produce alignments comparable to our gold-alignments. As reported in Table 2, awesome-align achieves an F1 score of only 0.1735 on our set of 500 gold aligned Tsez samples. Though we cannot assess the alignment quality of the automatic alignments we produced for Lezgi or Arapaho, we take this as an indication that our automatic alignments are of dubious accuracy, and this may adversely affect the performance of our approach. To address this possibility, we experiment with removing alignment from our approach entirely. We treat the cls-token embedding as a sentence-level representation of our translation input and use this in place of word-level embeddings. We then incorporate the embeddings as normal. We call this configuration CLS-Only.
6 Results and Discussion
We treat Tsez as our development language and perform a search over all translation incorporation strategies using the Tsez development set to determine the overall highest performing configuration, which we call TAMS. This configuration consists of Init-State in the encoder, Concat-Half in the decoder, and CLS-Concat as the CLS strategy. We additionally consider a variation we call TAMS-CLS, which is identical to TAMS except that it employs the CLS-Only strategy described in §5.3.1. Further details of our translation incorporation strategy search are included in §6.3.
6.1 Test Languages
We apply our final TAMS model configuration to the automatically aligned test sets in each of our three languages: Tsez (development language), Lezgi (in-family test language), and Arapaho (out-of-family test language). To simulate varying levels of data availability, we experiment with models trained on 100, 250, and 500 training samples in addition to experimenting with the full training set. For each of the train-limits, we report average metrics over five randomly chosen subsets of the full training set. There was a relatively high level of variance in the performance of our models, which should be considered when interpreting our results. The standard deviation of accuracy measures ranges between 1.11-2.33 across all settings. Full standard deviation metrics are reported in Appendix 11.
Results are shown in Table 1. In most cases, we see that edit distance, F1, and accuracy are roughly in agreement, so we focus on accuracy as our main evaluation metric. We find that on average our TAMS and/or TAMS-CLS outperforms the baseline in every train data setting and almost every metric. In general, performance gains are highest on the lower-resource settings, while on the higher-resource settings, performance improvements are slight if present at all. In the n=100 setting, TAMS outperforms the baseline by an average of 1.99 percentage points, suggesting that our model is most beneficial in truly low-data settings. We also see consistent gains on Arapaho, which indicates that with further work, our model could be useful for polysynthetic languages. This is particularly exciting considering the relative difficulty of segmenting polysynthetic languages.
Without assessments of alignment quality for each of our test sets, we cannot properly analyze the impact of poor alignments on TAMS, however we see that TAMS-CLS often outperforms TAMS. In fact, TAMS-CLS is often the highest performing model, which suggests that alignments are not strictly necessary to see performance gains from translation incorporation. We consider this a promising finding because of the relative difficulty of sourcing alignments.
6.2 Manual vs. Automatic Alignment
To directly compare the influence of automatic alignment on performance, we additionally train a TAMS model on gold-aligned Tsez data (Gold) and compare it to a TAMS model trained on equivalent data aligned with awesome-align. While we still report the average performance across 5 models trained on distinct training sets of 100 samples, the training sets we average here are distinct from the ones in §6.1) so performance metrics cannot be directly compared. To emphasize this point, we call the awesome-aligned setting Awesome-Gold. We report the performance of each model in Table 3. We find that TAMS-CLS is the highest performing model, suggesting that gold alignment data might not be necessary to see the best possible performance from TAMS.
It is possible that the granularity of information TAMS learns from alignments may in fact detract from its ability to generalize. This would be a positive result because expert alignment is costly, and without the need for good alignments, training a TAMS model becomes more feasible for a broader range of languages. However, further experimentation on more diverse languages is necessary to draw any conclusions to this end.
Tsez
Model Metrics Gold Awesome-Gold
Accuracy 32.68 31.12
TAMS F1 48.50 48.08
ED 705 714
Accuracy 33.96
TAMS-CLS F1 49.73
ED 705
Accuracy 33.48
Baseline F1 49.07
ED 727
Table 3: Manual vs. automatic alignment: Performance on Tsez’s gold-aligned test split (n=100).
6.3 Supplementary Analysis of Translation Incorporation Strategies